OpenStudy (anonymous):

Solve. x3 - 2x2 + x – 2 = 0

OpenStudy (anonymous):

factor by grouping is a good first step

OpenStudy (anonymous):

@SavannahWillett I agree with @completeidiot , you need to work on that factoring by grouping

OpenStudy (anonymous):

what are the first two terms?

OpenStudy (anonymous):

It should all be x^3 - 2x^2 + x -2 = 0

OpenStudy (anonymous):

yes but we're factoring it by grouping @SavannahWillett .... just like the one we just did

OpenStudy (anonymous):

Okay. So first we factor the first two, right? So they have x^2 in common?

OpenStudy (anonymous):

ok factor out the x^2 what's left?

OpenStudy (anonymous):

x-2

OpenStudy (anonymous):

and look at that, what are the last two terms?

OpenStudy (anonymous):

x - 2 xo

OpenStudy (anonymous):

So we have: \[X^2(X-2)+1(X-2)=0\]

OpenStudy (anonymous):

If you can factor by grouing, the last two terms will be a multiple of what you factor out... every time :) If they're not, then you can't factor by grouping

OpenStudy (anonymous):

Ohh, I get that now! What do we do after? Is that it?

OpenStudy (anonymous):

gotta factor it all the way now... factor out the (X-2)

OpenStudy (anonymous):

which x - 2?

OpenStudy (anonymous):

\[X^2(x-2)+1(X-2)\] (X-2) is the common factor

OpenStudy (anonymous):

Oh yeah! So its (x-2) (x^2-1) Right?

OpenStudy (anonymous):

close, look at your signs

OpenStudy (anonymous):

Crappy computer screen XD (x^2+1)(x-2)

OpenStudy (anonymous):

oh yeah blame it on the computer... lol ok so you have \[(X^2+1)(x-2)=0\] What are the roots?

OpenStudy (anonymous):

-1 and 2, yeah?

OpenStudy (anonymous):

2 is correct. if you plug in -1 to X^2 +1 what do you get?

OpenStudy (anonymous):

\[(-1)^2+1 = 2 \] not zero

OpenStudy (anonymous):

0

OpenStudy (anonymous):

oh Opps

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2 is the only real solution, the others are complex

OpenStudy (anonymous):

i and -i

OpenStudy (anonymous):

Awesome. I understand that now. a lot better than I did on the last problem XD